# Model Tells You What to Discard: Adaptive KV Cache Compression for LLMs

### v3

@@ Abstract
In this study, we introduce adaptive KV cache compression, a plug-and-play method that reduces the memory footprint of generative inference for Large Language Models (LLMs). Different from the conventional KV cache that retains key and value vectors for all context tokens, we conduct targeted profiling to discern the intrinsic structure of attention modules. Based on the recognized structure, we propose FastGen, which constructs the KV cache in an adaptive manner: evicting long-range contexts on attention heads emphasizing local contexts, discarding non-special tokens on attention heads centered on special tokens, and only employing the standard KV cache for attention heads that broadly attend to all tokens. Moreover, with the lightweight attention profiling used to guide the construction of the adaptive KV cache, FastGen can be deployed without resource-intensive fine-tuning or re-training. In our experiments across various asks, FastGen demonstrates substantial reduction on GPU memory consumption with negligible generation quality loss. We will release our code and the compatible CUDA kernel for reproducibility.
@@ 1 Introduction
[…]
2       ${\bm{K}}^{i}_{{\bm{C}}^{i}},{\bm{V}}^{i}_{{\bm{C}}^{i}}\leftarrow f({\bm{K}}^{i},{\bm{V}}^{i},{\bm{C}}^{i})$ $\hat{{\bm{K}}}^{i},\hat{{\bm{V}}}^{i}\leftarrow{\bm{K}}^{i}_{{\bm{C}}^{i}}{\bm{V}}^{i}_{{\bm{C}}^{i}}$
@@
1 ${\bm{z}}_{0}\leftarrow$ last prompt token for $j\in\{1,\cdots,\mbox{Max Generate Length}\}$  do
@@
3             ${\bm{K}}^{i},{\bm{Q}}^{i},{\bm{V}}^{i}\leftarrow H_{i}({\bm{z}}_{j-1},\hat{{\bm{K}}}^{i},\hat{{\bm{V}}}^{i})$ ${\bm{K}}^{i}_{{\bm{C}}^{i}},{\bm{V}}^{i}_{{\bm{C}}^{i}}\leftarrow f({\bm{K}}^{i},{\bm{V}}^{i},{\bm{C}}^{i})$ $\hat{{\bm{K}}}^{i},\hat{{\bm{V}}}^{i}\leftarrow{\bm{K}}^{i}_{{\bm{C}}_{i}},{\bm{V}}^{i}_{{\bm{C}}_{i}}$
@@ 2 Related Work
Underlying Structure of Attention. Inspired by the success of Transformer, extensive studies have been conducted to explore the underlying mechanism of different self-attention heads. Voita et al. (2019) analyzed the self-attention heads in BERT using LRF (Bach et al., 2015) and characterized them into interpretable roles, one of which is attending adjacent tokens all the time. Michel et al. (2019) demonstrated that heads in the same layer could have different impact on the performance while the importance of each head changes across tasks. Clark et al. (2019) and Kovaleva et al. (2019) identified such patterns as some heads primarily attend to separator tokens, adjacent tokens and a combination of these. While most previous studies mainly considered encoder models, FastGen is motivated by consistent patterns we have observed in decoder-only models. Like previous studies, FastGen also explores the structure of the attention mechanism to improve inference efficiency. But FastGan differs from previous studies by focusing on characterizing the KV cache of different attention heads.
@@ 3.1 Generative Inference of Autoregressive LLMs
Prompt Encoding. When an autoregressive transformer-based LLM generates the $i$-th token, the attention module needs to collect contextual information from all the preceding $i-1$ tokens, i.e., the key and value vectors (KV vectors) of these tokens. To circumvent redundant KV vector computations when generating succeeding tokens, all KV vectors are stored in the KV cache once they are generated.
@@ 3.2 FastGen
With these insights, we introduce FastGen: a dual-phase algorithm for crafting an adaptive KV cache. During the prompt encoding phase, model profiling is conducted to discern the behavior of various attention heads, so that we can choose the most appropriate compression strategy for each head. Then, in the token generation phase, instead of indiscriminately appending new KV vectors for each newly generated token, we manage the KV cache for each token based on its selected compression strategy.
@@ 3.3 Model Profiling
${\bm{C}}^{*}=\operatorname*{arg\,min}_{{\bm{C}}\in{\mathcal{C}}}\;\mbox{CacheMemoryCost}({\bm{C}})\;\;\mbox{ s.t. }\;\;|{\bm{A}}-\mbox{softmax}({\bm{Q}}{\bm{K}}_{{\bm{C}}}^{T})|\leq 1-T,$ (1)
@@
Intrinsically, our method assumes that the structure of the attention map is stable across different attention heads at different positions. So, it is sufficient to use only the encoded prompt to select a proper compression policy. It is worth noting that existing literature has provided the theoretical justification for using solely encoded prompts to capture attention structures for the full contexts (Zhang et al., 2023; Liu et al., 2023a). In our study, we also empirically verified this, as to be elaborated in Section 4.
@@ 4.1 Head Distinctive Attention Structure
Figure 3 shows that attention heads in different layers have vastly different structures. Specifically, for the initial and final layers, they have more attention heads assigned to the full KV cache, indicating attention heads in these layers are likely to attend to all tokens. Meanwhile, for middle layers, the attention map focuses on special tokens, indicating that most attention heads of these layers primarily attend to special tokens (i.e., the accumulated attention score on special tokens is higher than $0.95$ for these attention heads). Figure 1 shows the structure of different attention heads in the same layer. We see that attention structures differ across different layers and different heads.
@@ 4.2 Profile Tends to Be Consistent in One Sequence
The previous section demonstrates the great potential for constructing an adaptive KV cache in accordance with the structure of different attention heads. Here, we show that it is sufficient to leverage only the user-provided prompts and conduct one-shot model profiling, as outlined in Section 3.3. Specifically, we show that user-provided prompts share the same attention structure in the generation process.
@@ 5 Experiment
We conduct comprehensive experiments to demonstrate the effectiveness of FastGen on memory footprint reduction and generation quality preserving. First, we report the trade-off between memory reduction and end-to-end generation quality in Section 5.1, and discuss the compression ratio of FastGen in Section 5.2. To demonstrate the superiority of FastGen on real-world systems, we demonstrate the end-to-end latency change in Section 5.3 and the profiling overhead in Section 5.4. Finally, we present ablation studies and discussions in Section 5.5.
@@ 5.1 Trade-off between performance and memory reduction
We use standard generation tasks to evaluate Llama 1 and our fine-tuned Llama 1 models. For Llama 1, we choose four different tasks, including HumanEval (Chen et al., 2021), GSM8k (Cobbe et al., 2021), NQ (Kwiatkowski et al., 2019) and TQA (Kembhavi et al., 2017) to evaluate models’ abilities on different domains (code, math, question answering and reading comprehension). Note that in the four tasks, each testing sample is in a generative format, where answers are extracted after model generation finishes. This is crucial for a fair comparison on model’s generation quality. We evaluate the instruction finetuned LLaMa model on the instruction tuning benchmark AlpacaEval (Li et al., 2023), which consists of 805 question prompts from diverse domains.
@@
[…]
@@ 5.4 Profiling Cost
5.5 Ablations

### v4

@@ Abstract
In this study, we introduce adaptive KV cache compression, a plug-and-play method that reduces the memory footprint of generative inference for Large Language Models (LLMs). Different from the conventional KV cache that retains key and value vectors for all context tokens, we conduct targeted profiling to discern the intrinsic structure of attention modules. Based on the recognized structure, we propose FastGen, which constructs the KV cache in an adaptive manner: evicting long-range contexts on attention heads emphasizing local contexts, discarding non-special tokens on attention heads centered on special tokens, and only employing the standard KV cache for attention heads that broadly attend to all tokens. Moreover, with the lightweight attention profiling used to guide the construction of the adaptive KV cache, FastGen can be deployed without resource-intensive fine-tuning or re-training. In our experiments across various asks, FastGen demonstrates substantial reduction on GPU memory consumption with negligible generation quality loss.
@@ 1 Introduction
2       ${\bm{K}}^{i},{\bm{Q}}^{i},{\bm{V}}^{i}\leftarrow H_{i}(\mbox{Prompt})$
3       ${\bm{A}}^{i}\leftarrow\mbox{softmax}({{\bm{Q}}^{i}{\bm{K}}^{i}}^{T})$
4       ${\bm{C}}^{i}\leftarrow$ apply Equation 1 to ${\bm{A}}^{i}$
       /* Ci: optimal policy */
5       ${\bm{K}}^{i}_{{\bm{C}}^{i}},{\bm{V}}^{i}_{{\bm{C}}^{i}}\leftarrow f({\bm{K}}^{i},{\bm{V}}^{i},{\bm{C}}^{i})$
6       $\hat{{\bm{K}}}^{i},\hat{{\bm{V}}}^{i}\leftarrow{\bm{K}}^{i}_{{\bm{C}}^{i}}{\bm{V}}^{i}_{{\bm{C}}^{i}}$
7
@@ 4       ${\bm{C}}^{i}\leftarrow$ apply Equation 1 to ${\bm{A}}^{i}$
1 ${\bm{z}}_{0}\leftarrow$ last prompt token
2 for $j\in\{1,\cdots,\mbox{Max Generate Length}\}$  do
@@ 1 ${\bm{z}}_{0}\leftarrow$ last prompt token
4             ${\bm{K}}^{i},{\bm{Q}}^{i},{\bm{V}}^{i}\leftarrow H_{i}({\bm{z}}_{j-1},\hat{{\bm{K}}}^{i},\hat{{\bm{V}}}^{i})$
5             ${\bm{K}}^{i}_{{\bm{C}}^{i}},{\bm{V}}^{i}_{{\bm{C}}^{i}}\leftarrow f({\bm{K}}^{i},{\bm{V}}^{i},{\bm{C}}^{i})$
6             $\hat{{\bm{K}}}^{i},\hat{{\bm{V}}}^{i}\leftarrow{\bm{K}}^{i}_{{\bm{C}}_{i}},{\bm{V}}^{i}_{{\bm{C}}_{i}}$
7
@@ 2 Related Work
Underlying Structure of Attention. Inspired by the success of Transformer, extensive studies have been conducted to explore the underlying mechanism of different self-attention heads. Voita et al. (2019) analyzed the self-attention heads in BERT using LRF (Bach et al., 2015) and characterized them into interpretable roles, one of which is attending adjacent tokens all the time. Michel et al. (2019) demonstrated that heads in the same layer could have different impact on the performance while the importance of each head changes across tasks. Clark et al. (2019) and Kovaleva et al. (2019) identified such patterns as some heads primarily attend to separator tokens, adjacent tokens and a combination of these. While most previous studies mainly considered encoder models, FastGen is motivated by consistent patterns we have observed in decoder-only models. Like previous studies, FastGen also explores the structure of the attention mechanism to improve inference efficiency, but focusing on characterizing the KV cache of different attention heads.
@@ 3.1 Generative Inference of Autoregressive LLMs
Prompt Encoding. When an autoregressive transformer-based LLM generates the $i$-th token, the attention module needs to refer to all the preceding $i-1$ tokens, i.e., the key and value vectors (KV vectors) of these tokens. To circumvent redundant KV vector computations when generating succeeding tokens, all KV vectors are stored in the KV cache once they are generated.
@@ 3.2 FastGen
With these insights, we introduce FastGen: a dual-phase algorithm for crafting an adaptive KV cache. During the prompt encoding phase, model profiling is conducted to discern the behavior of various attention heads, so that we can choose the most appropriate compression strategy for each head. Then, in the token generation phase, instead of indiscriminately appending new KV vectors for each newly generated token, we manage the KV cache based on the selected compression strategies.
@@ 3.3 Model Profiling
$po{\bm{C}}^{*}=\operatorname*{arg\,min}_{{\bm{C}}\in{\mathcal{C}}}\;\mbox{CacheMemoryCost}({\bm{C}})\;\;\mbox{ s.t. }\;\;|{\bm{A}}-\mbox{softmax}({\bm{Q}}{\bm{K}}_{{\bm{C}}}^{T})|\leq 1-T,$ (1)
@@
Intrinsically, our method assumes that the structure of the attention map for a head is stable through the generation process. So, it is sufficient to use only the encoded prompt to select a proper compression policy. It is worth noting that existing literature has provided theoretical justifications for using solely encoded prompts to capture attention structures for the full contexts (Zhang et al., 2023; Liu et al., 2023a). In our study, we also empirically verified this, as to be elaborated in Section 4.
@@ 4.1 Head Distinctive Attention Structure
Figure 3 shows that attention heads in different layers have vastly different structures. Specifically, for the initial and final layers, they have more attention heads assigned to the full KV cache, indicating attention heads in these layers are likely to attend to all tokens. Meanwhile, for middle layers, the attention map focuses on special tokens, indicating that most attention heads of these layers primarily attend to special tokens (i.e., the accumulated attention score on special tokens is higher than $0.95$ for these attention heads). Figure 1 shows the structure of different attention heads in the same layer. We see that attention structures differ across different layers and heads.
@@ 4.2 Profile Tends to Be Consistent in One Sequence
The previous section demonstrates the great potential for constructing adaptive KV cache in accordance with the structure of different attention heads. Here, we show for each instance, it is sufficient to conduct one-shot model profiling, as outlined in Section 3.3. Specifically, for a given prompt, we show that the attention structure of each head remains consistent through the decoding process.
@@ 5 Experiment
We conduct comprehensive experiments to demonstrate the effectiveness of FastGen on memory footprint reduction and generation quality preserving. First, we report the trade-off between memory reduction and end-to-end generation quality in Section 5.1, and discuss the compression ratio of FastGen in Section 5.2. To demonstrate the superiority of FastGen on real-world systems, we demonstrate the end-to-end latency change in Section 5.3 and the profiling overhead in Section 5.4. Finally, we present ablation studies and discussions in Section A.1.
@@ 5.1 Trade-off between performance and memory reduction
We use standard generation tasks to evaluate Llama 1 and our fine-tuned Llama 1 models. For Llama 1, we choose 4 different tasks, including HumanEval (Chen et al., 2021), GSM8k (Cobbe et al., 2021), NQ (Kwiatkowski et al., 2019) and TQA (Kembhavi et al., 2017) to evaluate models’ abilities on different domains (code, math, question answering and reading comprehension). Note that in the four tasks, each testing sample is in a generative format, where answers are extracted after model generation finishes. This is crucial for a fair comparison on model’s generation quality. We evaluate the instruction finetuned LLaMa model on the instruction tuning benchmark AlpacaEval (Li et al., 2023), which consists of 805 question prompts from diverse domains.
@@
The evaluation of the Llama 1 model follows the default setting and evaluation metrics on each benchmark. We calculate F1 scores for GSM8k, NQ and TQA, and use the code execution Pass@1 rate for HumanEval. While evaluating an instruction-tuning model remains challenging, we follow previous work (Zhou et al., 2023; Touvron et al., 2023b) to use GPT4 as an evaluator for pair-wise comparison between two different model generations. For each prompt, we input the FastGen generation and the generation from the same model with Full KV Cache as a pair, and ask GPT4 to judge which one is better.We then calculate the win rate of FastGen over Full Cache. Hypothetically, the win rate of a lossless method should be around 50%. Aside from full-cache models, we also include non-adaptive KV cache methods for comparison. Specifically, we apply ${\bm{C}}_{\mbox{local}}$, ${\bm{C}}_{\mbox{frequent}}$, and ${\bm{C}}_{\mbox{local+frequent}}$ to all attention head without any adaptation, as baselines. It is worth mentioning that ${\bm{C}}_{\mbox{local+frequent}}$ is a very strong baseline as it is identical to the H2O method (Zhang et al., 2023) and the Scissorhands method (Liu et al., 2023a). We set $r_{l}=0.3$, $r_{f}=0.3$ in FastGen, and only change the recovery ratio $T$ to control the pruned KV cache ratio. For generation, we use nucleus sampling (Holtzman et al., 2019) with temperature T = 0.6, p = 0.9. Experiments are conducted on 8 NVIDIA A100 80GB GPUs.
@@ 5.4 Profiling Cost
6 Conclusion
Acknowledgments
Research was supported in part by US DARPA KAIROS Program No. FA8750-19-2-1004 and INCAS Program No. HR001121C0165, National Science Foundation IIS-19-56151, and the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of DARPA or the U.S. Government.
Appendix A Appendix
A.1 Ablations